# Ceci N’est Pas un Drone: Investigating the Impact of Design Representation on Design Decision Making When Using GenAI

### v1

@@ 3 Methodology
To test our hypotheses, we designed and conducted a within-subjects experiment across 3 populations to study whether and how different types of design representations, including design visualizations and numerical design performance data, affect engineers’ design choices and their ability to choose the optimal designs in a list of AI-generated design ideas.
@@ 3.1 Participants
In total, 160 participants were recruited, with 156 college students and 4 aerospace industry professionals. Out of the 156 recruited college students, 29 of them participated in study 1 as self-identified drone pilots and drone hobbyists from a U.S. research university. The other 127 of the college student participants in study 2 are from a sophomore-level engineering design course in the Department of Mechanical Engineering at the same U.S. research university. Participation was voluntary, and the participants were compensated with a \$10 Amazon gift card. Participants recruited from the design course are also compensated with course credits. The 4 aerospace industry participants in study 3 were recruited from a major aerospace company. The ethnicity, age, and gender of the participants did not affect the recruitment process. All participants were over the age of 18 when recruited.
@@
  • Among the 4 professional participants in study 3, all 4 of them identified as men. Also, 3 professional participants identified as White, and 1 identified as Native Hawaiian or Pacific Islander.
@@ 3.2 Study Design
The university ethics review board approves human-subjects research, and they approved this project. Each of the three studies consisted of one online survey. In the survey, participants individually consider three UAV (uncrewed aerial vehicles) design problems. The design problems are similar in terms of difficulty and design scope, but with different design requirements (e.g., low cost versus low maintenance). A detailed problem description and design requirements for each design problem are provided in Appendix A.
The UAV designs used in the survey are from the AircraftVerse dataset developed by SRI International and Southwest Research Institute cobb_aircraftverse_2023. The UAV designs in this dataset are created using a proprietary generative AI system developed by SRI International cobb_aircraftverse_2023 with real-world, off-the-shelf components with accurate physical properties. All the UAV designs used in this study are simulated for flight performance using a simulator developed for this work with the help of SRI International and Southwest Research Institute. We then documented the performance of each UAV design when conducting the tasks described in the design problems, enabling us to numerically determine the level of optimality of the design solutions presented to the participants.
The participants are asked to consider three design problems about UAVs with different design requirements. The three design problems are accompanied by three sets of design solutions. For each problem, they are tasked with selecting the best drone design and are also allowed to share the reasoning for their selection. The number of design solution options is 2, 8, and 16 for the three design problems, respectively. The 4-option case is used in training. The three sets of design solutions are the same for each experimental condition, but with different modalities, and are presented in random order. For each design problem, the three sets of design solutions are each presented by either 1) visual design rendering, 2) numerical design performance data with data visualization, or 3) visual design rendering plus numerical performance data and data visualization. The visual design rendering is presented with interactive 3D renderings of the UAV designs. The numerical design performance data is presented as tables, along with spider plots, which is a common practice in the aerospace industry. Figure 1 shows an example drone design using these two design modalities, as presented to the participants. The performance data available to the participants are max hover time, max travel distance, mass, max air speed, battery voltage, total cost, and max lift. Not all of the listed performance metrics are relevant to every design problem.
@@
The main identifying features and key performance metrics of the drone designs used in this study are shown in Tables LABEL:tab:drone_feature_1, LABEL:tab:drone_feature_2, and LABEL:tab:drone_feature_3 in Appendix B. We characterize each drone according to several identifying features; drone designs can be axis-symmetric or non-axisymmetric, planar or non-planar, and one-plane or off-plane. Some drones are axis-symmetric with respect to the center of the main body, with propellers aligned axisymmetrically to one another around the center of the main body of the drones on one horizontal plane. Non-axisymmetric designs are ones in which the propellers are not axisymmetric to each other around the center of the main body of the drones. Non-planar means that not all the propellers are on a single horizontal plane. Off-plane means that at least one propeller is on a different horizontal plane than the main body of the drone. We define designs that are axis-symmetric and one-plane as conventional designs, and designs that have atypical features (i.e., non-axisymmetric, non-planar, or off-plane) as unusual designs. Additionally, we introduce a measure Rod Length to Propeller Area Ratio which is the ratio between the total length of connecting rods of the drone and the total effective propeller area, with a unit of $m^{-1}$.
Key performance metrics include Maximum Thrust, Effective Lift, Hover Time, and Total Cost. Maximum Thrust is the maximum instant thrust that the drone can provide, based on our simulation, in Newtons ($N$). It is one of the most important features in evaluating the performance of the drone designs under the design requirements of the design problems, as it determines not only the amount of weight the drone can lift but also how quickly the drone can move. Effective Lift is the maximum weight ($kg$) that the drone can carry, determined by the maximum lift minus the drone’s own weight. Hover Time describes the maximum time that the drone can stay in the air and maintain the target position with minimal deviation. Total Cost is the total cost of the drone.
Neither Rod Length to Propeller Area Ratio nor Maximum Thrust is provided to the participants during the study. The former was introduced during analysis to support interpretation, while the latter was replaced with Max Lift to provide a more readily-interpretable metric of payload capacity.
The participant groups and design question orders of all three studies are shown in Table 1 below. Each study is detailed in the following subsections.
@@ 3.2.1 Study 1
For the first study, we targeted drone hobbyists and drone pilots at a U.S. research university. In total, 29 student drone hobbyists were recruited. In this study, the three design problems are presented in the order of 2-option, 8-option, and 16-option.
@@ 3.2.2 Study 2
In the first study (Study 1), we observed that the number of options could affect the participants’ choices, and that we see a drastically different trend with different design problems, especially those with 8 design options versus those with 2 or 16 options. We conducted the second study to identify whether the irregularity is due to the design problem itself or to the order in which the design problems were shown. We recruited 127 STEM students at the U.S. research university through an engineering design related course. The second study has two different survey design conditions. The first condition, Study 2A, presents the design problems in the order of 2-option, 8-option, and 16-option, identical to the first study. The second version of the survey, Study 2B, presents the design problems in the order of 2-option, 16-option, and 8-option. Both versions of the survey are identical in all aspects other than the order in which the design problems are shown. The design problems and the design options are the same as those used in the first study.
3.2.3 Study 3
Since the first two studies focused on the college student, we seek to examine potential differences between that novice population and experienced engineering professionals working in the aerospace industry. Thus, we ran a third study with participants who were working as aerospace engineers and designers in a large aerospace company. Since we did not find the order in which the design problems are presented to affect participants’ abilities to pick the better designs ($F(1,1135)=0.351,p=0.553$) in the second study, we kept the survey design identical to the first study, with the three design problems with 2, 8, and 16 options, in order.
@@ 3.3 Procedure
The participants in all three studies completed the study using an online survey tool (Qualtrics). After giving consent, the participants begin the study and are instructed to consider three design problems about UAVs with different design requirements. For each design problem, the design options are shown in three different types of design representation modalities, namely, visual rendering, numerical performance data with data visualization, and visual rendering with numerical performance data and data visualization. The participants are informed that the designs are automatically generated by an AI system and that we need their expertise to evaluate and determine the feasibility of these solutions. The participants are also informed that the designs are tested in an advanced simulator, but they still need to utilize their engineering experience and knowledge to evaluate the designs, considering real-world scenarios. The study then asks the participant to choose the most optimal drone design from the provided list of design options (see Appendix B) with respect to several design restrictions. For each design selection, the survey also asks the participants to briefly explain their choice in the form of a textual response.
@@ 3.4 Measurement and Data Analysis
In this section, we outline the metrics and analytical procedures used to evaluate participants’ decision-making across experimental conditions. The analysis incorporates different quantitative measures, capturing the variability and optimality of participants’ design selections as well as the thematic content of their written explanations. Together, these measures provide a comprehensive view of participants’ performance, enabling assessment of both the outcomes of their choices and the underlying reasoning processes.
@@ 3.4.1 Multiple Choice Question Response Entropy
We first calculated the Shannon entropy of the choices for each design problem. A smaller entropy value corresponds to greater order and less randomness in the participants’ choices, suggesting the participants are making similar decisions instead of random choices. The entropy is calculated as
@@ 3.4.2 Multiple Choice Question Response Accuracy
We also want to examine how different design modalities affect the participants’ ability to select better designs. In order to consistently select the best design, we employ the Pareto-TOPSIS (Technique for Order of Preference by Similarity to Ideal Solution) method hwang_multiple_1981; hwang_new_1993; hu_new_2023. Pareto-TOPSIS ranks candidate options based on their weighted distance to the ideal and the worst options, and is widely used in engineering design and HCI research as an objective means of establishing optimality hu_new_2023; alizadeh_data-driven_2019; souaille_interactive_2022; bertoni_iterative_2019; wang_enhanced_2021; chatterjee_probabilistic_2017. The option closest to the ideal and farthest from the worst is deemed the most optimal. In this work, we applied the same weight to all the design objectives in the design problems, namely, Hover Time, Maximum Thrust, Effective Lift, and Total Cost.
The Pareto-TOPSIS optimal solutions for problems with 2, 8, and 16 options are design 18393, design 20155, and design 20985, respectively. The Pareto-TOPSIS optimal designs (dominating Pareto optimal and best TOPSIS) are indicated with an asterisk in Tables LABEL:tab:drone_rendering_1-LABEL:tab:drone_feature_3 in Appendix B. We then calculated the accuracy of participants selecting the Pareto-TOPSIS optimal designs and performed ANOVA (Analysis of Variance) tests on the accuracies to unveil differences across experimental conditions.
@@ 3.4.3 Multiple Choice Question Response Change Across Modalities
We seek to examine how the participants picked different designs under different design representation modalities, and specifically how those selections changed between problems. We illustrated and analyzed the changes in participants’ design choices across design representation modalities with Sankey diagrams for each experiment.
3.4.4 Textual Response
Participants’ textual justification for their design choices is analyzed to better understand their decision making. We coded each participant’s textual responses based on the main driving feature(s) cited for selecting their chosen design. A code was only assigned if the participant explicitly mentioned the term or a near synonym (including different word forms).
We conclude 4 main driving features (non-mutually exclusive) for the visual rendering only condition, including symmetry, stability, reliability, and cost. These four factors are common contributing "visual inference" factors, where the participant is making their judgment based on the visualization of the design, including information inferred from the visualization. Please note that here cost is also inferred from the visual renderings as no numerical performance data is provided in this condition.
As for the numerical performance data only condition, we identified five main driving features (non-mutually exclusive), including hover time, lift, cost, stability, and reliability. Here, information about hover time, lift, and cost is provided with numerical data, and the participants also managed to infer information about stability, reliability of the design from the numerical data as well.
The main driving features (non-mutually exclusive) for the visual rendering with numerical performance data condition include symmetry, stability, reliability, hover time, lift, and cost. Here, the participants can infer design information from both the visual renderings and the provided performance metrics. Therefore, we categorize symmetry, stability, and reliability as visual inferences, as these factors are more likely to be inferred from the visual rendering. And, we categorize hover time, lift, and cost as performance metrics, as these factors are more likely to be obtained from the numerical performance data.
We then calculated and analyzed the frequency of those driving factors that appear in textual responses to find the main driving factors for participants’ decision-making.
4 Results
The results of each study will be shown separately in the following sub-sections, followed by the results of the analysis on the text responses. Each study subsection includes quantitative analysis on the multiple-choice question responses, including response entropy, accuracy, and design choice change across modalities.
4.1 Study 1 - Drone Hobbyists
The participants in Study 1 are drone hobbyists. The number of participants that have chosen each design is shown in Figures 2, 3, and 4, with detailed distribution available in Appendix C, Table 12, 13, and 14.
Figure 2: Study 1: Participants’ choice transition in Problem 1 - 2 options.
Figure 3: Study 1: Participants’ choice transition in Problem 2 - 8 options.
Figure 4: Study 1: Participants’ choice transition in Problem 3 - 16 options.
4.1.1 Multiple Choice Question Response Entropy
The entropy values of the design choices for each design problem for each experimental condition, along with the theoretical maximum and minimal entropy values, are shown in Figure 5 below. Overall, the entropy values are relatively small, indicating low randomness in the participants’ answers, suggesting the participants are making informed decisions based on the provided design information.
Figure 5: Study 1: Entropy of participants’ design choices
4.1.2 Multiple Choice Question Response Change Across Modalities
The consistency of participants’ design choices when presented with different design modalities is shown in Figure 6. Most participants’ design choices changed when the design modality changed from visual rendering to numerical performance data. Design modalities did affect participants’ decision-making here. However, there are no consistent changes found when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data. The transitions of participants’ design choice changes are shown in Figures 2, 3, and 4.
@@ 4.1.2 Multiple Choice Question Response Change Across Modalities
Participants’ selection of conventional design (axis-symmetric and one-plane) vs unusual design (non-axisymmetric, non-planar, or off-plane) is shown in Figure 7. The participants showed a strong preference for conventional designs when they were provided with only visual renderings. Also, fewer participants picked the unusual designs when visual renderings became available, when transitioning from numerical performance data to the mixed modality (both visual rendering and numerical performance data).
@@ 4.1.3 Multiple Choice Question Response Accuracy
The average results of drone hobbyist participants’ accuracy on selecting the optimal designs are shown in Figure 8, with error bars indicating a 95% confidence interval. The number of participants who have chosen the optimal designs is listed in Appendix C, Table 15.
Figure 8: Study 1: Participants’ accuracy on selecting the optimal designs. Error bars represent a 95% confidence interval. Participants are drone hobbyists.
We conducted ANOVA tests on the accuracy for each design problem. There are significant differences in participants’ accuracy in selecting the optimal designs when different design modalities are offered (Design problem 1 with 2 options: $F(2,84)=31.207$, $p<0.001$, $\eta_{p}^{2}=0.426$. Design problem 2 with 8 options: $F(2,84)=8.809$, $p<0.001$, $\eta_{p}^{2}=0.173$. Design problem 3 with 16 options: $F(2,84)=6.774$, $p=0.002$, $\eta_{p}^{2}=0.139$). The results indicate that the design modality does affect participants’ abilities to choose the better designs.
@@
Since different design problems result in different design choices, we looked into the differences caused by the design modality separately for each design problem. Specifically, we performed $t$-tests between experiment conditions, with a Bonferroni correction adjusted alpha value of 0.017, and the results are shown in Table 2. In more than half of the pairwise comparisons, there are significant differences in participants’ accuracies for selecting the optimal designs between the visual rendering and numerical data conditions, between the numerical data and mixed modality conditions, and between the visual rendering and mixed modality conditions. However, there are no significant differences found between numerical data and the mixed modality conditions for picking the optimal designs in the design problem with 16 design options. There are also no significant differences found between visual rendering and the mixed modality conditions for picking the optimal designs in both the design problem with 8 design options and the design problem with 16 design options.

### v2

@@ 3 Methodology
To test our hypotheses, we designed and conducted a within-subjects experiment across 2 populations to examine whether and how different types of design representations, including design visualizations and numerical design performance data, affect engineers’ design choices and their ability to select optimal designs from a list of AI-generated design ideas.
@@ 3.1 Participants
We recruited 156 college students from a U.S. research university. Of the 156 recruited participants, 29 participated in study 1 as self-identified drone pilots and drone hobbyists. The other 127 participants in study 2 are from a sophomore-level engineering design course in the Department of Mechanical Engineering. In this work, we specifically selected drone pilots and hobbyists for their experiential knowledge of drone performance and relevant domain knowledge. We anticipated that they might use this knowledge to augment their decisions. We selected engineering students as proxy participants for early-career engineers capable of understanding the design task and the goal of selecting optimal designs. All participants have the technical capabilities and engineering design knowledge to understand and perform the design tasks. It is a representative group to demonstrate the impact of different design modalities on design decision-making.
Participation was voluntary, and the participants were compensated with a \$10 Amazon gift card. Participants recruited from the design course are also compensated with course credits. Participants were only allowed to participate in one of the two studies. The ethnicity, age, and gender of the participants did not affect the recruitment process. All participants were over the age of 18 when recruited.
@@ 3.2 Study Design
The university ethics review board approves human-subjects research, and they approved this project. Each study used a Qualtrics online survey in which participants individually considered three UAV (uncrewed aerial vehicles) design problems. The design problems are similar in terms of difficulty and scope, but with different design requirements (e.g., low cost versus low maintenance). For example,
"The fire department is using UAVs to monitor wildfires. The UAV should have great hover time and carry at least 10 kg (22.0 lbs) of monitoring equipment. The UAV must rise to a designated height at a fast vertical lift speed. Without sacrificing hover time, more carrying capacity is desirable for carrying more equipment for better coverage. The UAV must also fly steadily and sustain cross-wind and other potential environmental hazards for safety reasons."
A detailed problem description and design requirements are provided in Appendix A.
The UAV designs used in the survey are from the AircraftVerse dataset developed by SRI International and Southwest Research Institute cobb_aircraftverse_2023. The UAV designs in this dataset are created using a proprietary generative AI system developed by SRI International cobb_aircraftverse_2023 with real-world, off-the-shelf components with accurate physical properties. All the UAV designs used in this study are simulated for flight performance using a simulator developed for this work, with support from SRI International and Southwest Research Institute. We then simulated and numerically determined the performance of each UAV design when conducting the tasks described in the design problems.
Each of the three design problems was accompanied by a set of candidate solutions. For each problem, participants were tasked with selecting the best drone design and were also allowed to share the reasoning for their selection. The number of design solution options is 2, 8, and 16 for the three design problems, respectively. A 4-option case is used for training. For each design problem, design solutions were presented in one of three representation modalities: 1) visual design rendering, 2) numerical design performance data with data visualization, or 3) visual design rendering plus numerical performance data and data visualization. Across modality conditions, the same sets of design solutions appeared, but in random order.
The visual design rendering is presented with interactive 3D renderings of the UAV designs. For the numerical design performance data with data visualization modality, we presented the performance data as tables along with spider plots. This choice was specifically advised by multiple aerospace professionals who identified the combination of tables and spider plots as a common practice in the aerospace industry. They also recommend against using aligned bar plots. Figure 1 shows an example drone design using these two design modalities, as presented to the participants. The performance data available to the participants are max hover time, max travel distance, mass, max air speed, battery voltage, total cost, and max lift. Not all of the listed performance metrics are relevant to every design problem.
@@
(c) Example design modality - visual rendering with numerical data and data visualization.
@@
The main identifying features and key performance metrics of the drone designs used in this study are shown in Tables LABEL:tab:drone_feature_1, LABEL:tab:drone_feature_2, and LABEL:tab:drone_feature_3 in Appendix B. We characterized each drone by several identifying features: drone designs can be axisymmetric or non-axisymmetric, planar or non-planar, and one-plane or off-plane. Some drones are axisymmetric with respect to the center of the main body, with propellers aligned axisymmetrically on a single horizontal plane. Non-axisymmetric designs lack such axisymmetry. Non-planar means that not all the propellers are on a single horizontal plane. Off-plane designs have at least one propeller on a different horizontal plane than the drone’s main body. We define axisymmetric and one-plane designs as conventional designs, and designs with any atypical features (i.e., non-axisymmetric, non-planar, or off-plane) as unusual designs.
Key performance metrics include Maximum Lift, Effective Lift, Hover Time, and Total Cost. Maximum Lift is the maximum lift force ($kg$) that the drone can provide. It is one of the most important features in evaluating the performance of the drone designs under the design requirements of the design problems, as it not only affects the amount of weight the drone can lift but also indicates how quickly the drone can move vertically. Effective Lift is the maximum weight ($kg$) that the drone can carry, determined by the maximum lift minus the drone’s own weight. Hover Time describes the maximum time that the drone can stay in the air and maintain the target position with minimal deviation. Total Cost is the total cost of the drone.
The participant groups and design question orders of the two studies are shown in Table 1 below. Each study is detailed in the following subsections.
@@ 3.2.1 Study 1
For the first study, we targeted drone hobbyists and drone pilots at a U.S. research university. In total, 29 student drone hobbyists were recruited. In this study, the three design problems were presented in the order of 2-option, 8-option, and 16-option.
@@ 3.2.2 Study 2
For the second study, we recruited 127 STEM students at the U.S. research university through an engineering design related course. The second study has two different survey design conditions. The first condition, Study 2A (63 participants), presents the design problems in the order of 2-option, 8-option, and 16-option, identical to the first study. The second version of the survey, Study 2B (64 participants), presents the design problems in the order of 2-option, 16-option, and 8-option. Both versions of the survey are identical in all aspects other than the order in which the design problems are shown. The design problems and the design options are the same as those used in the first study.  [^1] In Study 1, we initially observed a drastic difference in participants’ responses between the 8-option and the 16-option design problems, with our initial analysis method for design optimality. To determine whether the irregularity is significantly influenced by the order in which the design problems were presented, we conducted the second study with two different design problem orders. However, after data collection, we switched to the current Pareto-TOPSIS method for more objective and accurate optimality evaluation, and the irregularity is no longer observed.
@@ 3.3 Procedure
Participants completed the study using an online survey tool (Qualtrics). After giving consent, participants began the study and were instructed to consider three design problems about UAVs with different design requirements, with design options shown in different modalities. Participants are informed that the designs are automatically generated by an AI system and that we need their expertise to evaluate the feasibility of these solutions. The participants are also informed that the designs are tested in an advanced simulator, but they still need to utilize their engineering experience and knowledge to evaluate the designs, considering real-world scenarios. The study then asks the participant to choose the most optimal drone design from the provided list of options and briefly explain their choice in a textual response.
@@ 3.4 Measurement and Data Analysis
We analyzed participants’ responses using three quantitative measures: (1) response entropy, (2) response changes across modalities, and (3) response accuracy relative to optimal designs.
@@ 3.4.1 Multiple Choice Question Response Entropy
Before performing any analysis or statistical tests on the multiple-choice question responses, we first confirm that the participants are making informed decisions. To test if the question responses from the participants are rooted in the provided design information rather than randomness, we use Shannon entropy as a measurement of the level of information obtained by the participants. A smaller entropy value corresponds to greater order and less randomness in the participants’ choices, suggesting the participants are making similar decisions instead of random choices. The entropy is calculated as
@@
3.4.2 Multiple Choice Question Response Change Across Modalities
To understand the impact of different design modalities, we test whether the participants are making different design choices with different design modalities for the same design problem. Specifically, we seek to examine how the design choices changed between modalities. We illustrated and analyzed changes in participants’ design choices across design representation modalities, using figures showing the number of participants who changed design choices when modalities changed. A high number of participants who changed their design choices indicates that different design modalities influenced their decisions.
In addition, we illustrated and analyzed participants’ selection of conventional designs (axisymmetric and one-plane) versus unusual designs (non-axisymmetric, non-planar, or off-plane) with figures showing the number and percentage of participants who picked those designs. A higher number and percentage of participants selecting those designs reveal their preference for either conventional or unusual designs when presented with different design modalities.
@@ 3.4.3 Multiple Choice Question Response Accuracy
In addition to unveiling whether different design modalities affected participants’ design decisions, we examine how different design modalities affect the participants’ ability to select better designs.
In order to consistently select the best design, we employ the Pareto-TOPSIS (Technique for Order of Preference by Similarity to Ideal Solution) method hwang_multiple_1981; hwang_new_1993; hu_new_2023. Pareto-TOPSIS ranks candidate options based on their weighted distance to the ideal and the worst options, and is widely used in engineering design and HCI research as an objective means of establishing optimality hu_new_2023; alizadeh_data-driven_2019; souaille_interactive_2022; bertoni_iterative_2019; wang_enhanced_2021; chatterjee_probabilistic_2017. The option closest to the ideal and farthest from the worst is deemed the most optimal. In this work, we applied the same weight to all the design objectives in the design problems, namely, Hover Time, Maximum Lift, Effective Lift, and Total Cost.
The Pareto-TOPSIS optimal solutions for problems with 2, 8, and 16 options are design 18393, design 20155, and design 20985, respectively, and they are indicated with an asterisk in Tables LABEL:tab:drone_rendering_1-LABEL:tab:drone_feature_3 in Appendix B. We then calculated the accuracy of participants selecting the Pareto-TOPSIS optimal designs and performed ANOVA (Analysis of Variance) tests on the accuracies to unveil differences across experimental conditions. We also performed follow-up t-tests, with a Bonferroni correction adjusted alpha value, to compare participants’ accuracy with different design modalities as post hoc tests. Higher accuracy in selecting Pareto-TOPSIS optimal designs indicates a stronger ability to select better designs when participants are provided with that design modality.
4 Results
The results of each study will be shown separately in the following sub-sections. Each study subsection includes quantitative analyses of multiple-choice question responses, including response entropy, changes in design choice across modalities, and accuracy.
4.1 Study 1 - Drone Hobbyists
The participants in Study 1 are drone hobbyists. The number of participants that have chosen each design and the participants’ choice transitions are shown in Figures 10, 11, and 12 in Appendix C.
4.1.1 Multiple Choice Question Response Entropy
The entropy values of the design choices for each design problem are calculated (Design problem with 2 options: $H_{max}=1$, $H_{Visual}=0.401$, $H_{Data}=0.251$, $H_{Visual+Data}=0.678$. Design problem with 8 options: $H_{max}=3$, $H_{Visual}=1.273$, $H_{Data}=1.539$, $H_{Visual+Data}=1.500$. Design problem with 16 options: $H_{max}=4$, $H_{Visual}=1.401$, $H_{Data}=1.497$, $H_{Visual+Data}=1.551$). Overall, the entropy values are relatively small, indicating low randomness in the participants’ answers, suggesting the participants are making informed decisions based on the provided design information.
@@ 4.1.2 Multiple Choice Question Response Change Across Modalities
The consistency of participants’ design choices when presented with different design modalities is shown in Figure 2. Most participants’ design choices changed when the design modality changed from visual rendering to numerical performance data. Design modalities did affect participants’ decision-making here. However, there are no consistent changes found when the design modality changed from numerical performance data to the mixed modality with both visual rendering and numerical performance data. The transitions of participants’ design choice changes are shown in Figures 10, 11, and 12.
@@
Participants’ selection of conventional design (axisymmetric and one-plane) vs unusual design (non-axisymmetric, non-planar, or off-plane) is shown in Figure 3. The participants showed a strong preference for conventional designs when they were provided with only visual renderings. Also, fewer participants picked the unusual designs when visual renderings became available, when transitioning from numerical performance data to the mixed modality (both visual rendering and numerical performance data).
@@ 4.1.3 Multiple Choice Question Response Accuracy
The average results of drone hobbyist participants’ accuracy on selecting the optimal designs are shown in Figure 4, with error bars indicating a 95% confidence interval.
Figure 4: Study 1: Participants’ accuracy on selecting the optimal designs. Error bars represent a 95% confidence interval. p-value annotation legend: ns: 1.70e-02 < p <= 1.00e+00;    *: 1.00e-02 < p <= 1.70e-02;    **: 1.00e-03 < p <= 1.00e-02;    ***: 1.00e-04 < p <= 1.00e-03;    ****: p <= 1.00e-04. Participants are drone hobbyists.
We conducted ANOVA tests on the accuracy for each design problem. There are significant differences in participants’ accuracy in selecting the optimal designs when different design modalities are offered (Design problem with 2 options: $F(2,84)=31.207$, $p<0.001$, $\eta_{p}^{2}=0.426$. Design problem with 8 options: $F(2,84)=8.809$, $p<0.001$, $\eta_{p}^{2}=0.173$. Design problem with 16 options: $F(2,84)=6.774$, $p=0.002$, $\eta_{p}^{2}=0.139$). The results indicate that the design modality does affect participants’ abilities to choose the better designs.
@@
Further, we performed $t$-tests between experiment conditions, with a Bonferroni correction adjusted alpha value of 0.017, and the results are shown in Table 2. In more than half of the pairwise comparisons, there are significant differences in participants’ accuracies for selecting the optimal designs between the visual rendering and numerical data conditions, between the numerical data and mixed modality conditions, and between the visual rendering and mixed modality conditions. However, there are no significant differences found between numerical data and the mixed modality conditions for picking the optimal designs in the design problem with 16 design options. There are also no significant differences found between visual rendering and the mixed modality conditions for picking the optimal designs in both the design problem with 8 design options and the design problem with 16 design options.